OpenStudy (anonymous):

Algebra 1 question, medal and fan if able to help at all!

OpenStudy (anonymous):

Jackie has 3 square tablecloths with each side equal to k units. For each piece, she does something different to it according to each part below. Part C: Jackie cut off a strip of cloth from one side of the tablecloth and pasted the strip on an adjacent side of the tablecloth to increase its length by p units and decrease its width by p units. What will be the change in the area of the table cloth? Show your work. (3 points)

OpenStudy (anonymous):

@thomaster @Compassionate @whpalmer4 @amistre64

OpenStudy (anonymous):

I already finished most of the problem, I'm just stuck at the end: Area = (k+p)(k-p) = k(k) + k(-p) + p(k) + p(-p) =

OpenStudy (anonymous):

@eliassaab

OpenStudy (whpalmer4):

 Is that how you are interpreting this?